# R-Zero: Self-Evolving Reasoning LLM from Zero Data

### v3

@@ 1 Introduction
Self-evolving Large Language Models (LLMs) represent a promising frontier for advancing language intelligence. By autonomously generating, refining, and learning from their own experiences, these models provide a scalable pathway toward artificial superintelligence (Tao et al., 2024; Tan et al., 2024b). A critical requirement for training such self-evolving LLMs is access to large volumes of expertly curated tasks and labels, which serve as supervision signals for fine-tuning or reinforcement learning with verifiable rewards (RLVR) (Shao et al., 2024; DeepSeek-AI et al., 2025). However, relying on human annotators to create these tasks and labels is not only costly, labor-intensive, and difficult to scale, but also presents a fundamental bottleneck to advancing AI systems toward capabilities that could eventually surpass human intelligence (Su et al., 2025; Zhao et al., 2025a).
@@
To reduce dependence on human-curated data, self-generated and label-free methods have been proposed to eliminate the need for explicit supervision. In particular, label-free RL derives reward signals directly from the model’s own outputs, such as sequence-level confidence scores (Li et al., 2025a; Prabhudesai et al., 2025; Huang et al., 2025) and output entropy (Agarwal et al., 2025; Cheng et al., 2025). However, despite removing the need for explicit labels, label-free methods still relies on a pre-existing corpus of tasks, which limits its scalability in truly self-evolving settings. On the other side, self-challenging approaches train LLMs on tasks generated by the models themselves (Zhou et al., 2025a; Wang et al., 2025a; Zhao et al., 2025a), While promising, many of these methods rely on external code executors to ensure that the synthesized tasks are both feasible and verifiable. However, in domains that lack an explicit verification oracle, such as open-ended reasoning, ensuring the quality and correctness of self-generated data remains a significant challenge.
@@
In this paper, we propose R-Zero, a framework for training reasoning LLMs that can self-evolve from zero external data. In R-Zero, a single base model is initialized with two roles – a Challenger and a Solver that are independently optimized but co-evolve throughout the RL process. During co-evolving, the Challenger is rewarded for generating tasks targeted to be at the edge of Solver’s current abilities, while the Solver is rewarded for solving increasingly challenging tasks posed by the Challenger. Framework details are provided in Section 3, but briefly, in the Challenger training phase, the Challenger is trained via Group Relative Policy Optimization (GRPO) (Shao et al., 2024) to generate difficult questions. The reward signal is derived from the uncertainty for the frozen Solver, which is measured by the self-consistency of its multiple generated answers. In the Solver training phase, the Solver is fine-tuned with GRPO on a filtered set of these challenging questions generated by the now-frozen Challenger, using the pseudo-labels voted by itself. This entire process repeats, creating a self-evolving cycle that operates without any human intervention.
@@
Our experiments demonstrate that R-Zero is a model-agnostic framework, consistently and iteratively improving the reasoning abilities of different backbone LLMs. For example, Qwen3-4B-Base model’s average score on math benchmarks increased by a significant +6.49 points after three iterations of self-evolution. Moreover, the reasoning skills learned through our math-focused questions can generalize to complex general-domain tasks, with models trained using R-Zero showing significant improvements on general domain reasoning benchmarks like MMLU-Pro (Wang et al., 2024) and SuperGPQA (Du et al., 2025). Our further analysis finds that R-Zero can act as a mid-training method, as models first improved by our method achieve higher performance after fine-tuned on labeled data. In addition, we provide an in-depth analysis that validates our framework’s components, demonstrates its synergy with supervised fine-tuning, and characterizes the co-evolutionary dynamics to identify both strengths and limitations, offering insights for future research.
2 Preliminaries
2.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a reinforcement learning algorithm that fine-tunes a policy LLM $\pi_{\theta}$ without a separate, learned value function. Its core idea is to normalize rewards based on the performance of other responses generated from the same prompt.
For a given prompt $p$, a policy LLM $\pi_{\theta_{\text{old}}}$ generates a group of $G$ complete responses $\{x_{1},\dots,x_{G}\}$. Each response $x_{i}$ is evaluated to receive a single scalar reward $r_{i}$. The rewards across the group are then normalized using a z-score to compute a response-level advantage:
$\hat{A}_{i}=\frac{r_{i}-\operatorname{mean}(r_{1},\dots,r_{G})}{\operatorname{std}(r_{1},\dots,r_{G})+\varepsilon_{\text{norm}}},$
where $\varepsilon_{\text{norm}}$ is a small constant added for numerical stability.
Policy Update.
The policy is updated using a clipped surrogate objective, similar to PPO, to ensure stable training. The objective, regularized by a KL-divergence penalty to constrain policy drift, is:
$\mathcal{L}_{\text{GRPO}}(\theta)=-\frac{1}{G}\sum_{i=1}^{G}\min\!\Bigl(\tfrac{\pi_{\theta}(x_{i})}{\pi_{\theta_{\text{old}}}(x_{i})}\,\hat{A}_{i},\;\operatorname{clip}\!\bigl(\tfrac{\pi_{\theta}(x_{i})}{\pi_{\theta_{\text{old}}}(x_{i})},\,1-\epsilon,\,1+\epsilon\bigr)\,\hat{A}_{i}\Bigr)\;+\;\beta\,\operatorname{KL}\!\bigl(\pi_{\theta}\|\,\pi_{\theta_{\text{old}}}).$
Maximizing the negative of this loss encourages the policy to increase the probability of generating responses with positive relative advantages, while the KL term, controlled by $\beta$, limits divergence from the previous policy.
2.2 Reinforcement Learning with Verifiable Rewards
Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024) is a paradigm for fine-tuning models in domains where response quality can be deterministically verified. This approach relies on a rule-based verifier $v:\mathcal{X}\to\{0,1\}$ that assigns a binary reward to each generation $x_{i}$:
@@ 3.2 Challenger Training
The Challenger, $Q_{\theta}$, is an autoregressive language model trained to generate challenging questions. We train $Q_{\theta}$ using the GRPO algorithm detailed in Sec. 2. The core of this process lies in designing a reward function that accurately captures the desired properties of a “good” question. This final scalar reward, $r_{i}$, is then used in the GRPO advantage calculation. We focus on generating questions specifically within the domain of mathematics, as it provides a convenient and self-contained setting for our framework; the objective nature of mathematical answers allows for the straightforward generation of pseudo-labels via majority voting, without the need for external verification environments like code executors.
@@
This function incentivizes questions where the Solver is maximally uncertain (accuracy approaches 50%) (Sun et al., 2025). We provide a theoretical motivation for this reward function in Sec. 3.5.
@@
where $B$ is the batch size and $\lambda$ is a scaling factor. In our experiments, we set $\lambda=1$. The implementation details are shown in Appendix A.4.
@@ 3.4 Solver Training
$r_{j}=\begin{cases}1,&\text{if }x_{j}\text{ is identical to the pseudo-label }\tilde{y}_{i},\\ 0,&\text{otherwise}.\end{cases}$
@@
3.5 Theoretical Analysis
In this section, we provide a theoretical motivation for our uncertainty reward function, $r_{\text{uncertainty}}\propto 1-2|\hat{p}(x;S_{\phi})-\tfrac{1}{2}|$, which is maximized when the Solver’s success probability, $\hat{p}$, is 50%. Our analysis is grounded in recent work that formally establishes that the most efficient training occurs when a learner is exposed to tasks at the frontier of its capabilities (Shi et al., 2025a; Bae et al., 2025).
$\mathbb{D}_{KL}(S_{\phi}||S^{*})\geq\frac{\hat{p}(1-\hat{p})}{2\beta^{2}}$
where $\beta$ is the temperature parameter controlling entropy regularization. The right-hand side of the inequality, which is proportional to the reward variance, is maximized precisely when $\hat{p}=0.5$. Therefore, by designing the Challenger’s reward to incentivize questions that push the current Solver towards this point of maximum uncertainty, our framework is theoretically motivated to generate a maximally efficient curriculum in each iteration of the co-evolutionary process.
@@ 4 Experiments
4.1 Experimental Setting
4.1.1 Models
@@
We employ the Qwen3-4B-Base (Yang et al., 2025) and Qwen3-8B-Base models to assess the impact of scale within a single architectural family. Second, to ensure our approach is effective on a distinct lineage, we utilize the OctoThinker-3B and OctoThinker-8B models (Wang et al., 2025b).This choice is particularly relevant as Wang et al. (2025b) reported that applying RL training directly to Llama models yielded suboptimal results. As the OctoThinker series is continually trained from the Llama-3.1 models (Dubey et al., 2024), this comprehensive selection allows us to test our framework across different foundational architectures – Qwen vs. Llama.
4.1.2 Evaluation Benchmark
Mathematical Reasoning.
We use seven challenging benchmarks: AMC, Minerva (Lewkowycz et al., 2022), MATH-500 (Hendrycks et al., 2021b), GSM8K (Cobbe et al., 2021), Olympiad-Bench (He et al., 2024), AIME-2024, and AIME-2025. For these tasks, where answers can be complex, we employ GPT-4o as a programmatic judge to semantically verify the correctness of the final answer against the ground truth (Zhao et al., 2025c). For the difficult AMC and AIME benchmarks, we report the mean@32 metric. For all other math benchmarks, we report accuracy based on greedy decoding.
General Domain Reasoning.
To test for the generalization of reasoning ability, we evaluate on the following challenging benchmarks:
  • MMLU-Pro (Wang et al., 2024): An enhanced version of the MMLU (Hendrycks et al., 2021a) benchmark, featuring a more challenging suite of multi-task questions designed to provide a stricter evaluation of language model capabilities.
  • SuperGPQA (Du et al., 2025): A large-scale benchmark focused on graduate-level reasoning. It comprises questions across 285 distinct disciplines that have been verified as unsearchable on the web, thereby isolating true reasoning ability from simple knowledge recall.
  • BBEH (shoaa kazemi et al., 2025): This benchmark builds upon the foundation of BIG-Bench Hard (Suzgun et al., 2023) by incorporating a new selection of tasks specifically engineered to be more difficult, thus providing a more accurate measure of complex reasoning skills.
For this category, we follow the experimental setup, prompts, and evaluation codes from (Ma et al., 2025), reporting Exact Match (EM) accuracy obtained via greedy decoding.
4.1.3 Training Details
@@
Our entire framework is implemented based on the EasyR1 codebase (Zheng et al., 2025b). In each iteration of the R-Zero co-evolutionary loop, we follow a specific set of hyperparameters. The Challenger ($Q_{\theta}$) first generates a candidate pool of $N=8,000$ questions. To construct the training dataset for the Solver, these questions are filtered based on consistency. For each candidate question, we sample $m=10$ answers from the current Solver ($S_{\phi}$). A question is retained for the training set only if the number of answers matching the majority-vote pseudo-label is between 3 and 7, inclusive ($\delta=0.25$). This numerical range is consistent with the methodology used in previous research (Zhang and Zuo, 2025; Li et al., 2025b; Bercovich et al., 2025). When training the Challenger, the uncertainty reward $r(x;\phi)$ is calculated by sampling $m=10$ responses from the Solver. For the intra-batch repetition penalty, we set the clustering distance threshold to $\tau_{\text{BLEU}}=0.5$. Further implementation details and prompts can be found in Appendix A.
@@ 4.2 Results in Mathematical Reasoning
Table 1: Comprehensive results on mathematical reasoning benchmarks. We compare each base model against a Base Challenger baseline (where the Solver is trained on questions from an untrained Challenger) and our iterative method, R-Zero. The peak performance achieved during each model’s training process is highlighted in bold.
Model Name AVG AMC Minerva MATH GSM8K Olympiad AIME25 AIME24
@@
   Base Model 42.58 45.70 38.24 68.20 72.56 34.8 6.15 10.94
   Base Challenger 44.36 45.00 45.22 72.80 87.87 41.19 7.29 11.15
   R-Zero (Iter 1) 48.06 51.56 51.47 78.60 91.28 43.85 9.17 10.52
   R-Zero (Iter 2) 48.44 52.50 51.47 79.80 91.66 44.30 4.27 15.10
   R-Zero (Iter 3) 49.07 57.27 52.94 79.60 92.12 44.59 4.27 12.71
@@
   Base Model 49.18 51.95 50.00 74.40 89.08 42.74 16.67 13.85
   Base Challenger 51.87 60.70 57.72 81.60 92.56 46.44 13.44 10.62
   R-Zero (Iter 1) 53.39 61.56 59.93 82.00 93.71 48.00 14.17 14.37
   R-Zero (Iter 2) 53.84 61.56 59.93 82.00 93.93 48.30 17.60 13.54
   R-Zero (Iter 3) 54.69 61.67 60.66 82.00 94.09 48.89 19.17 16.35
@@
   Base Model 26.64 17.19 24.26 55.00 73.69 16.15 0.21 0.00
   Base Challenger 27.51 20.19 24.63 54.60 74.98 15.70 0.10 2.40
   R-Zero (Iter 1) 27.76 20.39 25.74 54.60 75.51 16.30 0.10 1.67
   R-Zero (Iter 2) 28.20 24.06 25.37 54.80 74.45 17.48 0.00 1.25
   R-Zero (Iter 3) 29.32 27.03 27.57 54.20 74.98 18.22 3.23 0.00
@@
   Base Model 36.41 32.11 41.91 65.20 86.96 26.52 1.56 0.62
   Base Challenger 36.98 29.30 42.28 66.20 88.10 27.56 1.04 4.38
   R-Zero (Iter 1) 37.80 32.97 45.22 65.60 86.96 28.44 1.98 3.44
   R-Zero (Iter 2) 38.23 32.58 48.53 67.20 87.11 27.26 0.00 4.90
   R-Zero (Iter 3) 38.52 34.03 48.22 68.80 87.19 27.56 0.42 3.44
@@
This improvement is progressive, with the results showing a clear trend of performance gains across iterations. For instance, the Qwen3-8B-Base model’s average score climbs from a base performance of 49.18 to 53.39 (Iter 1) and ultimately reaches 54.69 (Iter 3). A similar monotonic improvement is observed on OctoThinker-3B, which progresses from its base score of 26.64 to 29.32 after three iterations. This consistent growth underscores the benefits of the co-evolutionary dynamic, where the progressively more capable Solver learns from an increasingly challenging curriculum.
@@ 4.3 Results in General Reasoning
Previous work has demonstrated that training language models on reasoning-intensive domains, such as mathematics, can lead to improvements in general-domain capabilities (Huan et al., 2025). A key question, however, is whether this generalization effect still holds when the training curriculum is not human-labeled, but entirely self-generated through R-Zero.
As shown in Table 2, this transfer of skills is evident across all tested models. For instance, three iterations of our math-focused training improve the average general-domain score of Qwen3-8B-Base by +3.81 points and OctoThinker-3B by +3.65 points. This generalization also extends to the key performance patterns observed in the mathematical results, with progressive iterative gains. This confirms that our method does not merely teach domain-specific knowledge, but enhances the model’s underlying capabilities in a way that successfully generalizes across domains.

### v4

@@ 1 Introduction
Self-evolving Large Language Models (LLMs) represent a promising frontier for advancing language intelligence. By autonomously generating, refining, and learning from their own experiences, these models provide a scalable pathway toward artificial superintelligence (Tao2024ASO; tan2024large). A critical requirement for training such self-evolving LLMs is access to large volumes of expertly curated tasks and labels, which serve as supervision signals for fine-tuning or reinforcement learning with verifiable rewards (RLVR) (Shao2024DeepSeekMathPT; DeepSeekAI2025DeepSeekR1IR). However, relying on human annotators to create these tasks and labels is not only costly, labor-intensive, and difficult to scale, but also presents a fundamental bottleneck to advancing AI toward capabilities that could eventually surpass human intelligence (Su2025CrossingTR; Zhao2025AbsoluteZR).
@@
To reduce dependence on human-curated data, self-generated and label-free methods have been proposed to eliminate the need for explicit supervision. In particular, label-free RL derives reward signals directly from the model’s own outputs, such as sequence-level confidence scores (Li2025ConfidenceIA; Prabhudesai2025MaximizingCA; Huang2025EfficientTestTimeScalingSelfCalibration) and output entropy (Agarwal2025TheUE; Cheng2025ReasoningWE). However, despite removing the need for explicit labels, label-free methods still relies on a pre-existing corpus of tasks, which limits its scalability in truly self-evolving settings. On the other side, self-challenging approaches train LLMs on tasks generated by the models themselves (Zhou2025SelfChallengingLM; Wang2025CoEvolvingLC; Zhao2025AbsoluteZR). While promising, many of these methods rely on external code executors to ensure that the synthesized tasks are both feasible and verifiable. However, in domains that lack an explicit verification oracle, such as open-ended reasoning, ensuring the quality and correctness of self-generated data remains a significant challenge.
@@
In this paper, we propose R-Zero, a framework for training reasoning LLMs that can self-evolve from zero external data. In R-Zero, a single base model is initialized with two roles – a Challenger and a Solver that are independently optimized but co-evolve throughout the RL process. During co-evolving, the Challenger is rewarded for generating tasks targeted to be at the edge of Solver’s current abilities, while the Solver is rewarded for solving increasingly challenging tasks posed by the Challenger. Framework details are provided in Section 2, but briefly, in the Challenger training phase, the Challenger is trained via Group Relative Policy Optimization (GRPO) (Shao2024DeepSeekMathPT) to generate difficult questions. The reward signal is derived from the uncertainty for the frozen Solver, which is measured by the self-consistency of its multiple generated answers. In the Solver training phase, the Solver is fine-tuned with GRPO on a filtered set of these challenging questions generated by the now-frozen Challenger, using the pseudo-labels voted by itself. This entire process repeats, creating a self-evolving cycle that operates without any human intervention.
@@
Our experiments demonstrate that R-Zero is a model-agnostic framework, consistently and iteratively improving the reasoning abilities of different backbone LLMs. For example, Qwen3-4B-Base model’s average score on math benchmarks increased by a significant +6.49 points after three iterations of self-evolution. Moreover, the reasoning skills learned through our math-focused questions can generalize to complex general-domain tasks, with models trained using R-Zero showing significant improvements on general domain reasoning benchmarks like MMLU-Pro (Wang2024MMLUProAM) and SuperGPQA (Team2025SuperGPQASL). Our further analysis finds that R-Zero can act as a mid-training method, as models first improved by our method achieve higher performance after fine-tuned on labeled data. In addition, we provide an in-depth analysis that validates our framework’s components, demonstrates its synergy with supervised fine-tuning, and characterizes the co-evolutionary dynamics to identify both strengths and limitations, offering insights for future research.
@@ 2.2 Challenger Training
The Challenger, $Q_{\theta}$, is an autoregressive language model trained to generate challenging questions. We train $Q_{\theta}$ using the GRPO algorithm detailed in Sec. I. The core of this process lies in designing a reward function that accurately captures the desired properties of a “good” question. This final scalar reward, $r_{i}$, is then used in the GRPO. We focus on generating questions specifically within the domain of mathematics, as it provides a convenient and self-contained setting for our framework; the objective nature of mathematical answers allows for the straightforward generation of pseudo-labels via majority voting, without the need for external verification environments like code executors.
@@
This function incentivizes questions where the Solver is maximally uncertain (accuracy approaches 50%). We provide a theoretical motivation for this reward function in Appendix F.
@@
where $B$ is the batch size and $\lambda$ is a scaling factor. In our experiments, we set $\lambda=1$. The implementation details are shown in Appendix B.4.
@@ 2.4 Solver Training
$r_{j}=\begin{cases}1,&\text{if }x_{j}\text{ matches with the pseudo-label }\tilde{y}_{i},\\ 0,&\text{otherwise}.\end{cases}$
@@ 3 Experiments
3.1 Models and Training Details
@@
We employ the Qwen3-4B-Base (Yang2025Qwen3TR) and Qwen3-8B-Base models to assess the impact of scale within a single architectural family. Second, to ensure our approach is effective on a distinct lineage, we utilize the OctoThinker-3B and OctoThinker-8B models (wang2025octothinker).This choice is particularly relevant as wang2025octothinker reported that applying RL training directly to Llama models yielded suboptimal results. As the OctoThinker series is continually trained from the Llama-3.1 models (Dubey2024TheL3), this comprehensive selection allows us to test our framework across different foundational architectures – Qwen vs. Llama. We assess our framework on a comprehensive suite of benchmarks, with the evaluation benchmarks presented in Appendix C.
@@
Our entire framework is implemented based on the EasyR1 codebase (zheng2025easyr1). In each iteration of the R-Zero co-evolutionary loop, we follow a specific set of hyperparameters. The Challenger ($Q_{\theta}$) first generates a candidate pool of $N=8,000$ questions. To construct the training dataset for the Solver, these questions are filtered based on consistency. For each candidate question, we sample $m=10$ answers from the current Solver ($S_{\phi}$). A question is retained for the training set only if the number of answers matching the majority-vote pseudo-label is between 3 and 7, inclusive ($\delta=0.25$). This numerical range is consistent with the methodology used in previous research (Zhang2025GRPOLEADAD; li2025semanticallyawarerewardsopenendedr1; Bercovich2025LlamaNemotronER). When training the Challenger, the uncertainty reward $r(x;\phi)$ is calculated by sampling $m=10$ responses from the Solver. For the intra-batch repetition penalty, we set the clustering distance threshold to $\tau_{\text{BLEU}}=0.5$.
3.1.1 Evaluation Setting
The evaluation code is adopted from General-Reasoner (Ma2025GeneralReasonerAL). To ensure consistency, we reran the released evaluation code and reported the corresponding results. The reproduced results are well aligned with General-Reasoner and those in the Qwen-3 technical report (Yang2025Qwen3TR).
The results presented in all experimental tables are obtained after 45 training steps, while in the figures we report evaluations performed at checkpoints every 15 steps during solver training. All results are reported based on the held-out test sets, ensuring a fair comparison across baselines and reproduced methods. Further implementation details and prompts can be found in Appendix B.
@@ 3.2 Results in Mathematical Reasoning
Table 1: Comprehensive results on mathematical reasoning benchmarks. We compare each base model against a R-Zero ([Uncaptioned image] challenger) baseline (where the Solver is trained on questions from an untrained Challenger) and our method, R-Zero. The peak performance is highlighted in bold.
Model Name Avg. AMC Minerva MATH GSM8K Olympiad AIME25 AIME24
@@
   Base Model (w/o training) 42.57 45.70 38.24 68.20 87.79 41.04 10.30 6.7
   Absolute Zero 46.42 52.45 41.96 76.20 89.34 42.56 10.20 12.20
   R-Zero ([Uncaptioned image] challenger) 45.01 45.00 45.22 72.80 87.87 41.19 10.20 12.80
    R-Zero (our method) 49.93 57.27 52.94 79.60 92.12 44.59 9.60 13.40
@@
   Base Model (w/o training) 48.64 51.95 50.00 78.00 89.08 44.74 12.10 14.60
   Absolute Zero 52.68 57.89 57.90 76.60 92.00 47.80 18.20 18.40
   R-Zero ([Uncaptioned image] challenger) 52.10 60.70 57.72 81.60 92.56 46.44 11.60 14.10
    R-Zero (our method) 53.72 61.67 60.66 82.00 94.09 48.89 13.30 15.40
@@
   Base Model (w/o training) 26.64 17.19 24.26 55.00 73.69 16.15 0.21 0.00
   Absolute Zero 27.23 22.50 22.70 53.20 75.80 13.20 0.50 2.70
   R-Zero ([Uncaptioned image] challenger) 27.51 20.19 24.63 54.60 74.98 15.70 0.10 2.40
    R-Zero (our method) 29.32 27.03 27.57 54.20 74.98 18.22 3.23 0.00
@@
   Base Model (w/o training) 36.41 32.11 41.91 65.20 86.96 26.52 1.56 0.62
   Absolute Zero 36.60 32.50 44.90 62.80 87.00 25.60 3.30 0.10
   R-Zero ([Uncaptioned image] challenger) 36.98 29.30 42.28 66.20 88.10 27.56 1.04 4.38
    R-Zero (our method) 38.52 34.03 48.22 68.80 87.19 27.56 0.42 3.44
@@ 3.3 Results in General Reasoning
Previous work has demonstrated that training language models on reasoning-intensive domains, such as mathematics, can lead to improvements in general-domain capabilities (Huan2025DoesMR). A key question, however, is whether this generalization effect still holds when the training curriculum is not human-labeled, but entirely self-generated through R-Zero.
As shown in Table 2, this transfer of skills is evident across all tested models. For instance, three iterations of our math-focused training improve the average general-domain score of Qwen3-8B-Base by +5.13 points and OctoThinker-3B by +3.65 points. This generalization also extends to the key performance patterns observed in the mathematical results. This confirms that our method does not merely teach domain-specific knowledge, but enhances the model’s underlying capabilities in a way that successfully generalizes across domains.